# Self-intersection Number of Negative Curves on Fermat Surfaces

### v2

@@ 6.3. Resultants
$g=b_{0}v^{e}+b_{1}v^{e-1}+\cdots+b_{e},$
with $a_{0},b_{0}\neq 0$. The resultant $\mathrm{Res}(f,g)$ is the determinant
$\begin{vmatrix}a_{0}&a_{1}&a_{2}&\cdots&a_{d}&0&\cdots&0\\ 0&a_{0}&a_{1}&\cdots&a_{d-1}&a_{d}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots&\ddots&\ddots&\vdots\\ 0&0&0&a_{0}&\cdots&\cdots&a_{d-1}&a_{d}\\ b_{0}&b_{1}&b_{2}&\cdots&b_{e}&0&\cdots&0\\ 0&b_{0}&b_{1}&\cdots&b_{e-1}&b_{e}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots&\ddots&\ddots&\vdots\\ 0&0&0&b_{0}&\cdots&\cdots&b_{e-1}&b_{e}\end{vmatrix}$
which has $e$ rows of $a_{i}$ and $d$ rows of $b_{j}$.
Alternatively, if $\lambda_{1},\cdots,\lambda_{d}$ and $\mu_{1},\cdots,\mu_{e}$ are respectively the roots of $f,g$ in the algebraic closure of $K$, counted with multiplicities, then
$\mathrm{Res}(f,g)=a_{0}^{e}b_{0}^{d}\prod_{\begin{subarray}{c}1\leq i\leq d\\ 1\leq j\leq e\end{subarray}}(\lambda_{i}-\mu_{j}).$
For the polynomial $f$, the discriminant $\mathrm{Disc}(f)$ is defined as
$\mathrm{Disc}(f)=a_{0}^{2d-2}\prod_{i<j}(\lambda_{i}-\lambda_{j})^{2}.$
Proposition 6.4.
The following assertions are true:
  1. (i)
    $\mathrm{Disc}(f)=(-1)^{d(d-1)/2}a_{0}^{2d-2}\prod_{i\neq j}(\lambda_{i}-\lambda_{j})=\frac{(-1)^{d(d-1)/2}}{a_{0}}\mathrm{Res}\left(f,\frac{\partial f}{\partial v}\right).$
  2. (ii)
    $\mathrm{Res}(f,g)=(-1)^{de}\mathrm{Res}(g,f)$.
  3. (iii)
    If $d=0$, then $\mathrm{Res}(f,g)=a_{0}^{e}$.
  4. (iv)
    If $h$ is another polynomial in $K[v]$, then
    $\mathrm{Res}(fg,h)=\mathrm{Res}(f,h)\mathrm{Res}(g,h).$
  5. (v)
    If $a_{0}=1$, then for any $h\in K[v]$, we have
    $\mathrm{Res}(f,g)=\mathrm{Res}(f,g-fh).$
  6. (vi)
    If $f=f_{1}\cdots f_{k}$, then
    $\mathrm{Disc}(f)=\left(\prod_{i=1}^{k}\mathrm{Disc}(f_{i})\right)\left(\prod_{1\leq i<j\leq k}\mathrm{Res}(f_{i},f_{j})^{2}\right).$
6.5. Intersection multiplicities
Now we use the Puiseux series and resultants to study intersection multiplicities of plane curve germs. Suppose $F,G$ are two reduced curve germs in $(\mathbb{C}^{2},0)$, defined by
$F:\ f(u,v)=0,$
and
$G:\ g(u,v)=0.$
@@ 6.5. Intersection multiplicities
where $a_{i},b_{j}\in\mathbb{C}\{u\}$ satisfy $a_{i}(0)=b_{j}(0)=0$ for $1\leq i\leq d,1\leq j\leq e$. For convenience, we will also regard $a_{i},b_{j}$ as elements of $\mathbb{C}\langle\langle u\rangle\rangle$ by the inclusion $\mathbb{C}\{u\}\subset\mathbb{C}\langle\langle u\rangle\rangle$.
The intersection multiplicity of $F$ and $G$ at 00, denoted by $i_{0}(F,G)$, which is equal to the intersection multiplicity of $f$ and $g$ at 00, denoted by $i_{0}(f,g)$, is defined as
$i_{0}(F,G)=i_{0}(f,g)=\dim_{\mathbb{C}}\frac{\mathbb{C}\{u,v\}}{(f,g)}.$
The Milnor number of $F$ at 00, denoted by $\mu_{0}(F)$ or $\mu_{0}(f)$, is defined as the intersection multiplicity at 00 of $\frac{\partial f}{\partial u}$ and $\frac{\partial f}{\partial v}$:
$\mu_{0}(F)=i_{0}\left(\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)=\dim_{\mathbb{C}}\frac{\mathbb{C}\{u,v\}}{\left(\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)}.$
The following results are classical; the interested reader may consult [6, Chap.2] or [9, Chap.I, §3].
Proposition 6.6.
  1. (i)
    $i_{0}(f,g)<\infty$ if and only if $F$ and $G$ do not share a common irreducible component.
  2. (ii)
    If $h$ is another element in $\mathbb{C}\{u,v\},h(0,0)=0$, then $i_{0}(fg,h)=i_{0}(f,h)+i_{0}(g,h)$.
  3. (iii)
    If $F$ is irreducible and $s=s(u)$ is a $v$-root of $f$ such that $\nu(s)=d$, then
    $i_{0}(f,g)=do_{u}\left(g(u,s(u))\right)=\sum_{\alpha=1}^{d}o_{u}\left(g\left(u,s_{,\alpha}(u)\right)\right).$
  4. (iv)
    (Halphen’s formula) If $s_{1},\cdots,s_{d}$ and $s_{1}^{\prime},\cdots,s_{e}^{\prime}$ respectively are the $v$-roots of $f$ and $g$, then
    $i_{0}(f,g)=\sum_{\begin{subarray}{c}1\leq i\leq d\\ 1\leq j\leq e\end{subarray}}o_{u}(s_{i}-s_{j}^{\prime}).$
    As a corollary, we have
    $i_{0}(f,g)=o_{u}\left(\mathrm{Res}(f,g)\right).$
  5. (v)
    If $f=f_{1}\cdots f_{k}$, then
    $\mu_{0}(f)=\sum_{i=1}^{k}\mu_{0}(f_{i})+\sum_{i\neq j}i_{0}(f_{i},f_{j})-k+1.$
Now we establish the relationship between the Minor number and the discriminant.
@@
By [12, Lem.6.5.7, p.148], we have that the number
@@
Before presenting the proof of Proposition 4.4, we first recall our settings. We are given a reduced and irreducible curve $C$ on $X_{d}$, from which we obtain the following objects:
  1. (1)
    The branched Galois cover
    $\rho:\ X_{d}\to\mathbb{P}^{2},\quad[x,y,z,w]\mapsto[x,y,z].$
  2. (2)
    There exists a homogeneous polynomial $g(x,y,z)$ of degree $e$ such that
    $(g)=m(C_{1}+\cdots+C_{k}),$
    with $C_{i}^{2}=C^{2}$ for $i=1,\ldots,k$, and $k\geq 2$.
  3. (3)
    The subscheme $\mathscr{C}$ on $X_{d}$ defined by $g$, i.e.,
    $\mathscr{C}:\ g(x,y,z)=0.$
  4. (4)
    The curves $D$ and $E$ in $\mathbb{P}^{2}$ defined as
    $D:\ x^{d}+y^{d}+z^{d}=0\quad\mathrm{and}\quad E:\ g(x,y,z)=0.$
    In addition, $D$ and $E$ intersect at only finitely many points.
  5. (5)
    Fix an intersection point $P=[x_{0},y_{0},z_{0},w_{0}]$ of different irreducible components of $\mathscr{C}$. We have $w_{0}=0$. We also assume $x_{0}\neq 0$ and $z_{0}\neq 0$, and denote $p=[x_{0},y_{0},z_{0}]$.
  6. (6)
    Introduce the affine coordinates
    $u=\frac{y}{x}-\frac{y_{0}}{x_{0}},\qquad v=\frac{z}{x}-\frac{z_{0}}{x_{0}},\qquad\omega=\frac{w}{x},$
    and then the scheme $\mathscr{C}:g=0$ on $X_{d}$ is defined in $\mathbb{C}^{3}=\{x\neq 0\}\subset\mathbb{P}^{3}$ by
    $h(u,v)=0,\quad l(u,v)+\omega^{d}=0,$
    where
    $h(u,v)=g\left(1,u+\frac{y_{0}}{x_{0}},v+\frac{z_{0}}{x_{0}}\right)$
    and
  7. (7)
    Viewed as an element in $\mathbb{C}\{u,v\}$, the function $h$ is $v$-general of order $b$.
Now we are ready to give the proof of Proposition of 4.4. To begin with, we first decompose $h$ in $\mathbb{C}\{u,v\}$ into irreducible factors
@@
where $h_{1},\ldots,h_{r}$ are irreducible Weierstrass polynomial in $\mathbb{C}\{u\}[v]$ of degrees $b_{1},\ldots,b_{r}$ respectively, and $h_{0}$ is an invertible element in $\mathbb{C}\{u,v\}$. Recall that by assumption $h$ is $v$-general of order $b$, hence $b=b_{1}+\cdots+b_{r}$.
For $i\leq r$, let $s_{i}$ be a Puiseux series which is a root of $h_{i}$; without loss of generality, we assume $\nu(s_{i})=b_{i}$. Then, by Proposition 6.2 (iii) we have
$h_{i}=\prod_{\alpha=1}^{b_{i}}\left(v-s_{i,\alpha}(u)\right),\ i\leq r.$
(Recall that by our convention, if $s_{i}$ is written as
$s_{i}(u)=\sum a_{k}u^{\frac{k}{b_{i}}}$
then
$s_{i,\alpha}(u)=\sum e^{2\pi\sqrt{-1}\alpha k/b_{i}}a_{k}u^{\frac{k}{b_{i}}}$
for $\alpha=1,\ldots,b_{i}$.)
@@
$\phi_{i}(u,\omega)=\prod_{\alpha=1}^{b_{i}}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right),\ i\leq r,$
@@
    For $1\leq i<j\leq r$, $\phi_{i}$ and $\phi_{j}$ are coprime as elements in $\mathbb{C}\langle\langle u\rangle\rangle[\omega]$.
@@
Recall that $h_{i}$ and $h_{j}$ are coprime, so they do not have common roots in $\mathbb{C}\langle\langle u\rangle\rangle$. Therefore $s_{i,\alpha}(u)-s_{j,\beta}(u)\neq 0$.
@@
(B) To compute $\mu_{P}(\mathscr{C})$, we use Lemma 6.7. Then
@@
Hence, we need to calculate $o_{u}\left(\mathrm{Disc}(\Phi)\right)$. By Proposition 6.4 (vi), we have from $\Phi=\phi_{1}\cdots\phi_{r}$ the following equality
@@
$o_{u}\left(\mathrm{Disc}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right)\right)$ $=$ $(d-1)\left(l\left(u,s_{i,\alpha}(u)\right)\right)$
@@
Recall that
@@
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(l\left(u,s_{j,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\right)^{2d}.$
Recall that
$l\left(u,s_{j,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\simeq s_{j,\beta}(u)-s_{i,\alpha}(u);$
hence,
$\mathrm{Res}(\phi_{i},\phi_{j})^{2}$
@@
$=$ $(d-1)\sum_{i=1}^{r}i_{0}(h_{i},l)+do_{u}\left(\mathrm{Disc}(h)\right).$
Recall that $i_{0}(h,l)=i_{p}(E,D)$; and by Lemma 6.7, we have
@@
(C) Since $C_{1},\ldots,C_{k}$ are the irreducible components of $\mathscr{C}$, we have
@@
Summing over all possible intersection points and using Lemma 4.2, we obtain

### v3

@@ 5.2. Resultants, discriminants, and intersection multiplicities
G:g(u,v)=0.G:\ g(u,v)=0.
@@
where $a_{i},b_{j}\in\mathbb{C}\{u\}$ satisfy $a_{i}(0)=b_{j}(0)=0$ for $1\leq i\leq d,1\leq j\leq e$. For convenience, we will regard $a_{i},b_{j}$ as elements of $\mathbb{C}\langle\hskip-1.5pt\langle u\rangle\hskip-1.5pt\rangle$ by the inclusion $\mathbb{C}\{u\}\subset\mathbb{C}\langle\hskip-1.5pt\langle u\rangle\hskip-1.5pt\rangle$, and thus regard $f,g$ as polynomials in $\mathbb{C}\langle\hskip-1.5pt\langle u\rangle\hskip-1.5pt\rangle[v]$.
The resultant $\mathrm{Res}(f,g)$ is
$\mathrm{Res}(f,g)=\prod_{\begin{subarray}{c}1\leq i\leq d\\ 1\leq j\leq e\end{subarray}}(\lambda_{i}-\mu_{j}),$
where $\lambda_{1},\cdots,\lambda_{d}$ and $\mu_{1},\cdots,\mu_{e}$ are respectively the roots of $f,g$ in $\mathbb{C}\langle\hskip-1.5pt\langle u\rangle\hskip-1.5pt\rangle$. In addition, the discriminant $\mathrm{Disc}(f)$ is
$\mathrm{Disc}(f)=\prod_{i<j}(\lambda_{i}-\lambda_{j})^{2}.$
The intersection multiplicity of $F$ and $G$ at 0, denoted by $i_{0}(F,G)$, which is equal to the intersection multiplicity of $f$ and $g$ at 0, denoted by $i_{0}(f,g)$, is
$i_{0}(F,G)=i_{0}(f,g)=\dim_{\mathbb{C}}\frac{\mathbb{C}\{u,v\}}{(f,g)}=o_{u}\left(\mathrm{Res}(f,g)\right),$
where the last equality is the Halphen formula [6, Prop.2.6.3]. In addition, the Milnor number of $F$ at 0, denoted by $\mu_{0}(F)$ or $\mu_{0}(f)$, is
$\mu_{0}(F)=\mu_{0}(f)=i_{0}\left(\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)=\dim_{\mathbb{C}}\frac{\mathbb{C}\{u,v\}}{\left(\frac{\partial f}{\partial u},\frac{\partial f}{\partial v}\right)}.$
We establish the relationship between the Minor number and the discriminant.
@@
By [14, Lem.6.5.7], we have that the number
@@
Following the notations in Section 3, assume $h$ is $v$-general of order $b$, i.e.,
$h(0,v)=c_{b}v^{b}+c_{b+1}v^{b+1}+\cdots,$
for some constant $c_{b}\neq 0$.
We first decompose $h$ in $\mathbb{C}\{u,v\}$ into irreducible factors
@@
where $h_{1},\ldots,h_{r}$ are irreducible Weierstrass polynomial in $\mathbb{C}\{u\}[v]$ of degrees $b_{1},\ldots,b_{r}$ respectively, and $h_{0}$ is an invertible element in $\mathbb{C}\{u,v\}$, and $b=b_{1}+\cdots+b_{r}$.
For $1\leq i\leq r$, let $s_{i}$ be a Puiseux series which is a root of $h_{i}$. By [6, Cor.1.5.8, Thm.1.8.3], we have $\nu(s_{i})=b_{i}$ and
$h_{i}=\prod_{\alpha=1}^{b_{i}}\left(v-s_{i,\alpha}(u)\right),\ 1\leq i\leq r.$
@@
$\phi_{i}(u,\omega)=\prod_{\alpha=1}^{b_{i}}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right),\ 1\leq i\leq r,$
@@
    For $1\leq i<j\leq r$, $\phi_{i}$ and $\phi_{j}$ are coprime as elements in $\mathbb{C}\langle\hskip-1.5pt\langle u\rangle\hskip-1.5pt\rangle[\omega]$.
@@
Since $h_{i}$ and $h_{j}$ are coprime, they do not have common roots in $\mathbb{C}\langle\hskip-1.5pt\langle u\rangle\hskip-1.5pt\rangle$; therefore $s_{i,\alpha}(u)-s_{j,\beta}(u)\neq 0$.
@@
(B) To compute $\mu_{P}(\mathscr{C})$, we use Lemma 5.3 and get
@@
Hence, we need to calculate $o_{u}\left(\mathrm{Disc}(\Phi)\right)$.
From $\Phi=\phi_{1}\cdots\phi_{r}$, we have
$\mathrm{Disc}(\Phi)=\left(\prod_{i=1}^{r}\mathrm{Disc}(\phi_{i})\right)\left(\prod_{1\leq i<j\leq r}\mathrm{Res}(\phi_{i},\phi_{j})^{2}\right),$
and thus
@@
$o_{u}\left(\mathrm{Disc}\left(\omega^{d}+l\left(u,s_{i,\alpha}(u)\right)\right)\right)$ $=$ $(d-1)o_{u}\left(l\left(u,s_{i,\alpha}(u)\right)\right)$
@@
In the proof of (A), we have shown that
@@
$\simeq$ $\prod_{\alpha=1}^{b_{i}}\prod_{\beta=1}^{b_{j}}\left(l\left(u,s_{j,\beta}(u)\right)-l\left(u,s_{i,\alpha}(u)\right)\right)^{2d}$
@@
$=$ $(d-1)\sum_{i=1}^{r}i_{0}(h_{i},l)+do_{u}\left(\mathrm{Disc}(h)\right)$
$=$ $(d-1)i_{0}(h,l)+do_{u}\left(\mathrm{Disc}(h)\right).$
Note that $i_{0}(h,l)=i_{p}(E,D)$; and by Lemma 5.3, we have
@@
(C) Since $C_{1},\ldots,C_{k}$ are the irreducible components of $\mathscr{C}$, we have from [14, Thm.6.5.1] that
@@
Summing over all possible intersection points $P\in C_{i}\cap C_{j}$ with $i\neq j$ and using Lemma 4.2, we obtain
@@
Data Availability: This manuscript has no associated data. All results are derived through rigorous mathematical proofs.
Funding Declaration: This work was previously supported by the Chinese Academy of Sciences (CAS) under Grant No. YSBR-001.